# Distributed Discrete-time Dynamic Outer Approximation of the Intersection of Ellipsoids

### v1

@@ VI Application: distributed Kalman filter
The previous result motivates estimating (4) in a distributed way:
$\mathbf{Q}_{i}[k]\hat{\mathbf{x}}_{i}[k]=\frac{1}{{\textsf{N}}}\sum_{j\in\mathcal{I}}\left(\frac{\lambda_{\mathbf{P}}^{j}[k]}{1-\sum_{l\in\mathcal{N}_{j}}\lambda_{l}^{j}[k]}\right)\mathbf{P}_{j}[k]^{-1}\bar{\mathbf{x}}_{j}[k].$ (5)
@@
Consider the same setting as in Proposition 1. Moreover, let ${\hat{\theta}}=1$. Then, the update rule in eq. (4) leads to
$\lim_{k\to\infty}\|\hat{\mathbf{x}}_{i}[k]-\hat{\mathbf{x}}[k]\|=0$
@@
First, let
$\lambda^{*}_{i}=\frac{1}{{\textsf{N}}}\left(\frac{\lambda_{\mathbf{P}}^{i}}{1-\sum_{j\in\mathcal{N}_{i}}\lambda_{j}^{i}}\right),$
where $\lambda^{i}_{\mathbf{P}}$ and $\lambda_{j}^{i}$ come from (2) in equilibrium at agent $i\in\mathcal{I}$. Then,
@@
and $\sum_{j=1}^{\textsf{N}}\lambda_{j}^{*}\leq 1$ with $\lambda_{j}^{*}\geq 0$. Moreover, if Assumption 2 is satisfied, (6) is complied with equality. Now, note that Lemma 5 implies that in equilibrium all agents are in consensus at $\mathbf{Q}_{i}[k]=\mathbf{Q}^{*}$ with constant $\mathbf{Q}^{*}=\mathbf{Q}^{*}[K]$. Moreover, due to the constraints in (2):
$\mathbf{Q}_{i}[k]\kern-2.84544pt=\kern-2.84544pt\mathbf{Q}^{*}\kern-2.84544pt\preceq\kern-2.84544pt\lambda^{i}_{\mathbf{P}}\mathbf{P}_{i}[k]^{-1}\kern-2.84544pt+\kern-5.69046pt\sum_{j\in\mathcal{N}_{i}}\kern-2.84544pt\lambda_{j}^{i}\mathbf{Q}^{*}\rightarrow\mathbf{Q}^{*}\kern-2.84544pt\left(\kern-2.84544pt1-\kern-5.69046pt\sum_{j\in\mathcal{N}_{i}}\kern-2.84544pt\lambda_{j}^{i}\kern-2.84544pt\right)\kern-2.84544pt\preceq\kern-2.84544pt\lambda^{i}_{\mathbf{P}}\mathbf{P}_{i}[k]^{-1}.$
Now, dividing by $1-\sum_{j\in\mathcal{N}_{i}}\lambda_{j}^{i}$ and applying a summation over all agents in the previous relation: ${\textsf{N}}\mathbf{Q}^{*}\preceq\sum_{i=1}^{\textsf{N}}\left(\frac{\lambda_{\mathbf{P}}^{i}}{1-\sum_{j\in\mathcal{N}_{i}}\lambda_{j}^{i}}\right)\mathbf{P}_{i}[k]^{-1}$ which reduces to (6) by the definition of $\lambda_{i}^{*}$. Now, note that $0\leq\lambda_{\mathbf{P}}^{i}\leq 1-\sum_{j\in\mathcal{N}_{i}}\lambda_{j}^{i}$ implying $\lambda_{i}^{*}\geq 0$ and also that,
$\lambda_{i}^{*}\leq\frac{1}{{\textsf{N}}}\left(\frac{1-\sum_{j\in\mathcal{N}_{i}}\lambda_{j}^{i}}{1-\sum_{j\in\mathcal{N}_{i}}\lambda_{j}^{i}}\right)=\frac{1}{{\textsf{N}}}$
Hence, $\sum_{j=1}^{\textsf{N}}\lambda_{j}^{*}\leq 1$. In addition, if Assumption 2 is satisfied, the all previous inequalities are interchanged with equalities.
@@
The results of Proposition 1 and 2 lead to a novel distributed Kalman filter with the following four steps: at instant $k$ and given $\hat{\mathbf{x}}_{i}[k-1],\hat{\mathbf{P}}_{i}[k-1]$, each node (i) predicts $\bar{\mathbf{x}}_{i}[k],\bar{\mathbf{P}}_{i}[k]$ using the known linear stochastic dynamics of the target system $\mathbf{x}[k]=\mathbf{A}\mathbf{x}[k-1]+\mathbf{w}[k]$, with $\mathbf{w}[k]$ a zero-mean Gaussian noise with covariance $\mathbf{W}$; (ii) exchanges of $\bar{\mathbf{x}}_{i}[k],\bar{\mathbf{P}}_{i}[k],\mathbf{Q}_{i}[k-1]$ with neighbors $i\in\mathcal{N}_{i}$; (iii) uses Algorithm 1 to obtain $\mathbf{Q}_{i}[k]$, which is employed as the new covariance matrix to be updated; (iv) updates the predictions using (4) and the Kalman filter update equations using the measurements $\mathbf{y}_{i}[k]=\mathbf{H}_{i}\mathbf{x}_{i}[k-1]+\mathbf{v}_{i}[k]$, with $\mathbf{v}_{i}[k]$ a zero-mean Gaussian noise with covariance $\mathbf{V}_{i}$, obtaining $\hat{\mathbf{x}}_{i}[k],\hat{\mathbf{P}}_{i}[k]$. We refer to [42, 6] for the explicit expressions for the prediction and correction steps of distributed Kalman filtering.
@@ VII-A Constant inputs
First, we evaluate Algorithm 1 in the static case, i.e., $\mathbf{P}_{i}[k]=\mathbf{P}_{i}$ $\forall i\in\mathcal{I}$. We generate a random connected graph $\mathcal{G}$ of $\mathsf{N}=6$ nodes, leading to the following edge set: =$\mathcal{E}=\{(1,2),(1,3),(1,5),(2,4),(2,5),(2,6)\}$. Input matrices are of dimension $n=2$, and they are initialized as $\mathbf{P}_{i}=\mathbf{L}_{i}^{\top}\mathbf{L}_{i}$, where the elements of each $\mathbf{L}_{i}$ are randomly generated using a uniform distribution between $-1$ and $1$. The resulting ellipses are characterized by the following matrices:
@@
Figure 4: Illustrative example of Algorithm 1 applied to a static problem. At each time step, the error between the eigenvalues of the global optimum matrix and the local estimates $|q^{i}_{j}[k]-q^{*}_{j}|$ decreases until the error becomes zero in finite time (in this case, $K=3$). The top left pannel shows the initial ellipsoids at each node and the optimal solution from the original centralized Löwner-John method. The top right pannel shows, for node $1$, its estimate after using Algorithm 1 at instant $k=1$ and the ellipsoids exchanged with its neighbors. The bottom right pannel depicts the estimate at node $1$ and instant $k=2$ after using again Algorithm 1, depicting that node $1$ already recovers the desired optimum $\mathbf{Q}^{*}$. The bottom left pannel depicts the evolution of the error with time for all the nodes.
@@ VII-C Application example: distributed Kalman filtering
Fig. 6 reports the average of the mean square error across nodes and simulation runs for the two distributed Kalman filters:
$\text{MSE}[k]=\frac{1}{100}\sum_{s=1}^{100}\frac{1}{\mathsf{N}}\sum_{i=1}^{\mathsf{N}}||\bar{\mathbf{x}}^{s}_{i}[k]-\mathbf{x}^{s}[k]||^{2}.$
Our proposed algorithm surpasses CO-DKF by exploiting a better approximation of the true fused covariance matrix under known correlations: while our approach tracks the global minimum across all the nodes, CO-DKF only computes the outer approximation of the intersection of the neighboring ellipsoids. The difference between both approximations is greater for larger and sparser networks.

### v2

@@ VI Application: distributed Kalman filter
Proposition 1 exploits consistency to guarantee that the estimates at each node can be fused when cross-correlations are unknown, either for static of dynamic input matrices. The previous result motivates estimating (4) in a distributed way:
$\mathbf{Q}_{i}[k]\hat{\mathbf{x}}_{i}[k]=\frac{1}{{\textsf{N}}}\sum_{j\in\mathcal{I}}\left(\frac{\lambda_{\mathbf{P}}^{j}[k]}{1-\sum_{l\in\mathcal{N}_{j}}\lambda_{l}^{j}[k]}\right)\mathbf{P}_{j}[k]^{-1}\bar{\mathbf{x}}_{j}[k]{,}$ (5)
@@
Consider the same setting as in Proposition 1. Moreover, let ${\bar{\theta}}=1$. Then, the update rule in eq. (4) leads to $\lim_{k\to\infty}\|\hat{\mathbf{x}}_{i}[k]-\hat{\mathbf{x}}[k]\|=0.$
@@
First, let $\lambda^{*}_{i}=\frac{1}{{\textsf{N}}}\left(\frac{\lambda_{\mathbf{P}}^{i}}{1-\sum_{j\in\mathcal{N}_{i}}\lambda_{j}^{i}}\right),$ where $\lambda^{i}_{\mathbf{P}}$ and $\lambda_{j}^{i}$ come from (2) in equilibrium at agent $i\in\mathcal{I}$. Then,
@@
and $\sum_{j=1}^{\textsf{N}}\lambda_{j}^{*}\leq 1$ with $\lambda_{j}^{*}\geq 0$. Moreover, if Assumption 2 is satisfied, (6) is complied with equality. Now, note that Lemma 4 implies that in equilibrium all agents are in consensus at $\mathbf{Q}_{i}[k]=\mathbf{Q}^{*}$ with constant $\mathbf{Q}^{*}=\mathbf{Q}^{*}[K]$. Moreover, due to the constraints in (2): $\mathbf{Q}_{i}[k]\kern-2.84544pt=\mathbf{Q}^{*}\preceq\lambda^{i}_{\mathbf{P}}\mathbf{P}_{i}[k]^{-1}+\sum_{j\in\mathcal{N}_{i}}\kern-2.84544pt\lambda_{j}^{i}\mathbf{Q}^{*}$ implies $\mathbf{Q}^{*}\kern-2.84544pt\left(-\sum_{j\in\mathcal{N}_{i}}\kern-2.84544pt\lambda_{j}^{i}\kern-2.84544pt\right)\kern-2.84544pt\preceq\kern-2.84544pt\lambda^{i}_{\mathbf{P}}\mathbf{P}_{i}[k]^{-1}.$ Now, dividing by $1-\sum_{j\in\mathcal{N}_{i}}\lambda_{j}^{i}$ and applying a summation over all agents in the previous relation: ${\textsf{N}}\mathbf{Q}^{*}\preceq\sum_{i=1}^{\textsf{N}}\left(\frac{\lambda_{\mathbf{P}}^{i}}{1-\sum_{j\in\mathcal{N}_{i}}\lambda_{j}^{i}}\right)\mathbf{P}_{i}[k]^{-1}$ which reduces to (6) by the definition of $\lambda_{i}^{*}$. Now, note that $0\leq\lambda_{\mathbf{P}}^{i}\leq 1-\sum_{j\in\mathcal{N}_{i}}\lambda_{j}^{i}$ implying $\lambda_{i}^{*}\geq 0$ and as a result that $\lambda_{i}^{*}\leq 1/{\textsf{N}}$. Hence, $\sum_{j=1}^{\textsf{N}}\lambda_{j}^{*}\leq 1$. In addition, if Assumption 2 is satisfied, the all previous inequalities are interchanged with equalities.
@@
The results of Proposition 1 and 2 lead to a novel distributed Kalman filter with the following four steps: at instant $k$ and given $\hat{\mathbf{x}}_{i}[k-1],\hat{\mathbf{P}}_{i}[k-1]$, each node (i) predicts $\bar{\mathbf{x}}_{i}[k],\bar{\mathbf{P}}_{i}[k]$ using the known linear stochastic dynamics of the target system $\mathbf{x}[k]=\mathbf{A}\mathbf{x}[k-1]+\mathbf{w}[k]$, with $\mathbf{w}[k]$ a zero-mean Gaussian noise with covariance $\mathbf{W}$; (ii) exchanges of $\bar{\mathbf{x}}_{i}[k],\bar{\mathbf{P}}_{i}[k],\mathbf{Q}_{i}[k-1]$ with neighbors $i\in\mathcal{N}_{i}$; (iii) uses Algorithm 1 to obtain $\mathbf{Q}_{i}[k]$, which is employed as the new covariance matrix to be updated; (iv) updates the predictions using (4) and the Kalman filter update equations using the measurements $\mathbf{y}_{i}[k]=\mathbf{H}_{i}\mathbf{x}_{i}[k-1]+\mathbf{v}_{i}[k]$, with $\mathbf{v}_{i}[k]$ a zero-mean Gaussian noise with covariance $\mathbf{V}_{i}$, obtaining $\hat{\mathbf{x}}_{i}[k],\hat{\mathbf{P}}_{i}[k]$. We refer to [36, 9] for the explicit expressions for the prediction and correction steps of distributed Kalman filtering. It is interesting to remark that, even though Eq. (4) is similar to a batch version of CI, it is not since it incorporates information from all nodes through $\mathbf{Q}^{*}[k],\lambda_{j}^{*}$, which are computed using Algorithm 1. Hence, while existing methods are limited to local CI, our method can calculate a global CI once Algorithm 1 has converged.
@@ VII-A Constant inputs
First, we evaluate Algorithm 1 in the static case, i.e., $\mathbf{P}_{i}[k]=\mathbf{P}_{i}$ $\forall i\in\mathcal{I}$. We generate a random connected graph $\mathcal{G}$ of $\mathsf{N}=6$ nodes, leading to the following edge set: =${\mathcal{F}}=\{(1,2),(1,3),(1,5),(2,4),(2,5),(2,6)\}$. Input matrices are of dimension $n=2$, and they are initialized as $\mathbf{P}_{i}=\mathbf{L}_{i}^{\top}\mathbf{L}_{i}$, where the elements of each $\mathbf{L}_{i}$ are randomly generated using a uniform distribution between $-1$ and $1$. The resulting ellipses are characterized by the following matrices:
@@
Figure 4: Illustrative example of Algorithm 1 applied to a static problem. At each time step, the error between the eigenvalues of the global optimum matrix and the local estimates $|q^{i}_{j}[k]-q^{*}_{j}|$ decreases until the error becomes zero in finite time (in this case, $K=3$). The top left pannel shows the initial ellipsoids at each node and the optimal solution from the original centralized Löwner-John method. The top right pannel shows, for node $1$, its estimate after using Algorithm 1 at $k=1$ and the ellipsoids exchanged with its neighbors. The bottom right pannel depicts the estimate at node $1$ and $k=2$ after using again Algorithm 1, depicting that node $1$ already recovers the desired $\mathbf{Q}^{*}$. The bottom left pannel depicts the evolution of the error with time for all the nodes.
@@ VII-C Application example: distributed Kalman filtering
Fig. 6 reports the average of the mean square error across nodes and simulation runs for the two distributed Kalman filters: $\text{MSE}[k]=\frac{1}{100}\sum_{s=1}^{100}\frac{1}{\mathsf{N}}\sum_{i=1}^{\mathsf{N}}||\bar{\mathbf{x}}^{s}_{i}[k]-\mathbf{x}^{s}[k]||^{2}.$ Our proposed algorithm surpasses CO-DKF by exploiting a better approximation of the true fused covariance matrix under known correlations: while our approach tracks the global minimum across all the nodes, CO-DKF only computes the outer approximation of the intersection of the neighboring ellipsoids. The difference between both approximations is greater for larger and sparser networks.